Can someone help me on this problem?
I need to solve the following word problem and need to show the equation you use for the solution.
Martina leaves home at 9 a.m., bicycling at a rate of 24 mi/h. Two hours later, John leaves driving at the rate of 48 mi/h. At what time will John catch up with Martina?
Thanks,
Ashley

Answer by stanbon(75887)

Martina leaves home at 9 a.m., bicycling at a rate of 24 mi/h. Two hours later, John leaves driving at the rate of 48 mi/h. At what time will John catch up with Martina?
----------------
Martina DATA:
rate=24 mph ; time = x hrs. ; distance = rt = 24x miles
----------------------
John DATA:
rate=48 mph ; time= x-2 hrs. ; distance = rt = 48(x-2)=48x-96 miles
--------
EQUATION:
distance = distance
24x = 48x-96
x=4 hrs
John meets Martina at 9:00AM + 4 hr= 1:00PM
Cheers,
Stan H.